The rate of change of the value of an investment, S, with respect to time, t ≥ 0, is given by dS dt = e(rt/100)(5r), where r is the annual interest rate (assumed constant) and the principal of the investment is S(0) = 500.The rate of change of the value of an investment, S, with respect to time, t ≥ 0, is given by


1
Expert's answer
2020-06-10T18:23:18-0400
"\begin{aligned}

\frac{dS}{d t}=5r \ e^{\frac{r t}{100}}, s(0)=500 \quad(\text { given })\\[2em]
\therefore d S=5 r\ e^{\frac{r}{100} t} d t \quad \text { (integrate both sides) }

\end{aligned}\\
\begin{aligned} \\
\int_{0}^{t} d S&=\int_{0}^{t} 5 r \ e^{\frac{r t}{100}} d t\\[1 em]
S\bigg| _{0}^{t}& =5 r\left(\frac{100}{r}\right) e^{\frac{r}{100} t}\bigg| _{0}^{t} \\[1 em]
S(t)-S(0) &=500\left(e^{\frac{r t}{100}}-1\right)\\[1 em]
 S(t) &=500\left(e^{\frac{r t}{100}}-1\right)+S(0) \\[1 em]
&=500\left(e^{\frac{r t}{100}}-1\right)+500 \\[1 em]
&=500\ e^{\frac{r t}{100}}-500+500 \\[1 em]
 \therefore S (t)&=500\ e^{\frac{r t}{100}}\\[1 em]
 \end{aligned}\\
\text{To verify the solution we find the derivative of }\\[1 em]

\frac{dS}{d t}=500 (\frac{r}{100}) \ e^{\frac{r t}{100}}=5r \ e^{\frac{r t}{100}}"